### v1

# Large Row-Constrained Supersaturated Designs for High-throughput Screening

Abstract
High-throughput screening, in which multiwell plates are used to test large numbers of compounds against specific targets, is widely used across many areas of the biological sciences and most prominently in drug discovery. We propose a statistically principled approach to these screening experiments, using the machinery of supersaturated designs and the Lasso. To accommodate limitations on the number of biological entities that can be applied to a single microplate well, we present a new class of row-constrained supersaturated designs. We develop a computational procedure to construct these designs, provide some initial lower bounds on the average squared off-diagonal values of their main-effects information matrix, and study the impact of the constraint on design quality. We also show via simulation that the proposed constrained row screening method is statistically superior to existing methods and demonstrate the use of the new methodology on a real drug-discovery system.
Keywords: experimental design; biological screening; regularization
1 Introduction
High-throughput screening (HTS) of chemical or biological compounds is used extensively across a range of critical scientific enterprises. This type of screening is common in basic functional genomics studies and synthetic biology applications (Zhou et al., 2014; Zeng et al., 2020; Sarnaik et al., 2020), but arguably its highest impact application is drug discovery (Perola, 2010; Blay et al., 2020; Soheilmoghaddam et al., 2021; Aldewachi et al., 2021). The literature suggests that nearly one-third of clinical drug candidates emerge from high-throughput screening campaigns (Brown and Boström, 2018). These screens were a centerpiece in the National Cancer Institute’s Cancer Moonshot program (Wilson et al., 2020). But as liquid handling, miniaturization, and other screening technologies have improved, the number of compounds to explore has increased as well (Elkin et al., 2015; Volochnyuk et al., 2019). This, in turn, highlights the need to improve the experimental and statistical aspects of these screens. In the current work, we present a scalable statistical methodology, including a novel experimental design approach, to make high-throughput screening more flexible, cost-effective, and statistically efficient for detecting promising compounds, compared to existing methods. We do this by proposing a new type of supersaturated design which assigns a number of compounds to each well. This design is then analyzed using regularized regression. These new designs are characterized by a constraint which governs the number of compounds which can be included in each well. Using the unconditional $E(s^{2})$ supersaturated design criterion, we present a computationally efficient heuristic to construct these designs, and study them by developing lower bounds on the criterion and investigating how design quality changes as a function of the constraint. We also give evidence of the effectiveness of our approach via simulation. Finally, we demonstrate the methodology in real experimental screens used to identify an inhibitor of a metallo-$\beta$-lactamase. The prevalence of such enzymes is of great concern because they exist in bacteria that cause nosocomial infections, including in secondary bacterial infections contributing to an estimated 50% of COVID-19 deaths (Mojica et al., 2022; Chedid et al., 2021). Searching for compounds that inhibit these enzymes’ activity is an important part of the arsenal for addressing antibiotic resistance (CDC, 2019).
HTS utilizes plates with varying well capacities (e.g., 96, 384, or 1536). Each well contains cells or enzymes and is subjected to application-specific conditions. Most commonly, a single chemical compound is assigned to a well, a biological or chemical process then occurs, and a measurement is taken. By using many plates, huge compound libraries can be studied. Compounds that produce meaningful measurements are then validated by more focused experimentation. The process is called screening because the goal is to detect, among thousands of candidates, the few compounds that exhibit the most promise. Though experimental design and statistical analysis have been promoted in HTS, and sophisticated methods of normalization and bias-correction have been developed (Altman, 2005; Birmingham et al., 2009; Malo et al., 2010; Dragiev et al., 2011; Zhang, 2011; Caraus et al., 2015), the basic experimental design approach most often used is simplistic: apply each treatment to one well (Malo et al., 2006; Kainkaryam and Woolf, 2009) (or, as recommended in the literature, in triplicate; see Malo et al., 2010) and then use a z-score to determine if the result is a hit. As shown in Section 4, this approach’s advantage of simplicity comes at the cost of reduced ability to detect hits. We note the recent work using machine/deep learning to preselect the “best” compounds (e.g., Stokes et al., 2020; Cheng et al., 2024), and these approaches have become popular for drug discovery (Mayr et al., 2018; Camacho et al., 2018). Our approach is complementary to these efforts, in the sense that it might be used to create training sets for HTS machine/deep learning models.
We are not the first to propose pooling multiple compounds in each well; in fact, versions of pooling have been used successfully (see Kainkaryam and Woolf, 2009). The most aggressive version of the existing strategies is called non-adaptive pooling (Thierry-Mieg, 2006; Thierry-Mieg and Bailly, 2008; Wilson et al., 2020; Kainkaryam and Woolf, 2009), which purportedly can screen, with a high degree of accuracy, large numbers of compounds with relatively few pools. The most relevant approach uses so-called Shifted Transversal Designs (STDs) within a larger procedure called poolHiTS. This method requires that the response for each well can be identified as either a “hit” or “inert”, and also that the user specifies the expected number of hits and the expected error rate. STDs construct pools to have as little overlap between them as possible, and they use information about which pools trigger hits to back out which compounds are likely hits and which are inert. The chief advantage of STDs is that they are fast to construct, but they have several important drawbacks described in more detail in Section 4.1.2. Most basically, poolHiTS treats neither the design nor the analysis in a statistical way. Our approach, on the other hand, frames the problem as a statistically designed experiment, which generates data posited to arise from a statistical model. This allows us to use statistical estimation procedures and thereby control false positive rates with high true positive rates. Overall, we will show that our procedure is more flexible with better statistical performance than poolHiTS.
Our approach uses a supersaturated design (SSD) to guide the placement of multiple compounds in each well, where $k$ compounds are investigated using $n$ wells (with $n\leq k$) while controlling false positive rates and increasing true positive rates. SSDs—experiments for which there are more factors than runs—were first described over sixty years ago (Satterthwaite, 1959; Booth and Cox, 1962). Motivated by the desire to experiment on complex high-cost systems, SSDs have seen extensive development over the last thirty years (e.g., Lin, 1993; Wu, 1993; Nguyen, 1996; Cheng, 1997; Cheng and Tang, 2001; Butler et al., 2001; Bulutoglu and Cheng, 2004; Georgiou, 2014; Jones and Majumdar, 2014; Weese et al., 2021). In our setting, an SSD is a two-level design composed of $k$ compounds (factors; absence with level $-1$; presence with level $+1$), and $n$ wells (runs) with $k\geq n$. The SSD provides a set of instructions regarding which compounds are included in which well. The design is then analyzed assuming the main-effects statistical model
$\mathbf{y}=\beta_{0}\mathbf{1}+X\bm{\beta}+\bm{\epsilon},$ (1)
where $X$ is the $n\times k$ SSD design matrix with elements $x_{ij}=\pm 1$, $\bm{\beta}=(\beta_{1},\dots,\beta_{k})^{T}$ and $\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}I)$. Because $k\geq n$, ordinary least squares cannot be used to estimate factor effects. Although forward selection approaches (Westfall et al., 1998; Abraham et al., 1999; Kelly III and Voelkel, 2000) were initially considered, regularization methods are now preferred for the analysis of SSDs. The Gauss-Dantzig selector (Candes and Tao, 2007) has been shown to be effective (Phoa et al., 2009; Marley and Woods, 2010; Weese et al., 2015), but the Lasso (Tibshirani, 1996) is closely related to the Dantzig selector (Meinshausen et al., 2007; Lounici, 2008; James et al., 2009; Asif and Romberg, 2010) and achieves similar performance (see Draguljić et al., 2014). Additionally, Stallrich et al. (2023) has recently developed an optimal screening design framework based on the Lasso. Thus, in our work we use the Lasso, which selects estimates according to
$(\hat{\beta}_{0},\hat{\bm{\beta}})=\operatorname*{arg\,min}_{\beta_{0},\bm{\beta}}\frac{1}{2n}||\bm{y}-\beta_{0}\bm{1}-X\bm{\beta}||_{2}^{2}+\lambda||\bm{\beta}||_{1},$ (2)
where $||\bm{\beta}||_{\ell}=(\sum_{j=0}^{k}|\beta_{j}|^{\ell})^{1/\ell}$ and $\lambda>0$.
A critical aspect of the proposed methodology is constructing SSDs to use as maps guiding the assignment of compounds to wells. SSDs are classically constructed using the $E(s^{2})$ criterion (Booth and Cox, 1962; Nguyen, 1996; Cheng, 1997; Bulutoglu and Cheng, 2004) to create designs with columns balanced between $-1$’s and $+1$’s, as close to mutually orthogonal as possible. We follow work that relaxes the balance requirement (Marley and Woods, 2010; Jones and Majumdar, 2014; Weese et al., 2015), leading to the unconditional $E(s^{2})$ criterion. Mathematically, let $L=[\mathbf{1},X]$, and $S=L^{{}^{\prime}}L$ with elements $s_{ij}$ and $O_{S}=\sum_{i<j}s_{ij}^{2}$. Then the unconditional $E(s^{2})$ criterion is
$UE(s^{2})=\frac{O_{S}}{k(k+1)},$ (3)
the average of the squared off-diagonal elements of $S$. In many cases $UE(s^{2})$-optimal designs can be constructed directly from known Hadamard matrices (Jones and Majumdar, 2014), but in our work we construct the designs algorithmically; see Section 2.1.
Our approach, which we call constrained row screening, or CRowS, breaks new ground in several important ways. First, it provides a statistically principled and substantially improved solution to an important problem in biology. We will demonstrate empirically in Section 4 that our approach has much better statistical properties, in terms of true and false positive rates, than existing methods. Secondly, the constraints inherent to the HTS application necessitate innovation in design construction. Specifically, though there is literature suggesting that multiple compounds can be applied to individual wells (e.g., Elkin et al., 2015; Kainkaryam and Woolf, 2008, 2009), it is also clear that in drug screening settings, and by extension high-throughput screening more broadly, there are limitations regarding the number of compounds that can be added to each well. These may be due to the size of the well, the minimum concentration required in order for a compound to be detectable, or the maximum ionic strength of the solution beyond which the compounds may precipitate. We call these limitations the row constraints, denoted by $c$, because it limits the number of $+1$’s in each row of the design. Thus, we have developed a new coordinate-exchange algorithm (Meyer and Nachtsheim, 1995) that aims to minimize $UE(s^{2})$ while respecting row constraints, and we study these new CRowS designs both theoretically and empirically as a function of $c$. Third, our work provides a significant, broad-ranging application for SSDs, which has been lacking in the literature. Despite their considerable promise (Weese et al., 2021), we only know of a handful of scattered articles that report on their use (Carpinteiro et al., 2004; Dejaegher and Vander Heyden, 2008; Jridi et al., 2015; Zarkadas and Besseris, 2023). Our work changes that, by applying these designs to a large class of biological/chemical screens and demonstrating their usefulness in real experiments. Finally, SSD research has focused largely on designs with at most dozens of factors; in contrast, CRowS designs will necessarily be larger in order to apply them to HTS. The designs we investigate in this paper have more than one hundred factors, though the proof-of-concept real experiments we report on in Section 5 are smaller. Since HTS studies routinely run into the tens of thousands and even much larger (Elkin et al., 2015), the size of the SSDs applied to this area will be limited not by the number of compounds studied, but by, in large part, logistical constraints related to the liquid handling automation (see Section 5 for a discussion) as well as computational limitations in the construction of large designs. We note that Eskridge et al. (2004) study large SSDs, and even given a nod to HTS as an application, but they don’t provide designs for $n>24$ nor do they consider row constraints or provide a real example.
The remainder of the paper is structured as follows. In Section 2, we describe row-constrained supersaturated designs in more detail, along with the algorithm we use to construct them. Section 3 provides some initial lower bounds on the $UE(s^{2})$ values for these CRowS designs, as well as an empirical study that sheds light on their behavior as a function of the severity of the row constraint. In Section 4 we use a set of extensive simulations to show the effectiveness of our proposed procedure compared to alternative methods. In Section 5 we provide more detail on the metallo-$\beta$-lactamase system, and describe the results of several screens as a real-world proof-of-concept of CRowS. We provide a discussion and conclusion in Section 6.
2 Row-Constrained Supersaturated Designs
In this section we describe the new class of supersaturated designs that have emerged from this work, along with the heuristic optimization procedure we have developed to construct them. As explained in Section 1, $UE(s^{2})$-optimal designs minimize the average of the squared off-diagonal elements of information matrix $S$. In the present setting, we propose a generalization of $UE(s^{2})$-optimal designs that impose a constraint on the number of $+1$’s in each row. Thus, using notation from Section 1, we say that an $n\times k$ design $X$ is row-constrained $UE(s^{2})$-optimal if $n\leq k$, $[X]_{ij}=x_{ij}$, and the design solves the following optimization problem:
min $UE(s^{2})$ (4)
s.t. $x_{ij}\in\{-1,1\}\;\forall i,j$ (5)
$\sum_{j=1}^{k}x_{ij}\leq 2c-k\;\forall i$ (6)
where $UE(s^{2})$ is defined in (3) and $c$ is the maximum number of compounds/well. As mentioned in the Introduction, we call the proposed row-constrained $UE(s^{2})$-optimal designs constrained row screening, or CRowS, designs.
2.1 Construction Algorithm
In order to construct approximately optimal row-constrained $UE(s^{2})$ designs, we use the coordinate exchange (CEx) heuristic (Meyer and Nachtsheim, 1995), along with some computational devices to make the procedure faster. Let $n_{1i}$ be the number of $+1$’s in design row $x_{i}$, so that $n_{1i}\leq c\;\forall i$. The algorithm is described in Algorithm 1. For the designs in this paper, we ran Algorithm 1 with 100 random starts and chose the design with the best objective function value.
Algorithm 1 Row-Constrained Design Construction
1:Randomly generate an initial design with $n_{1i}\leq c$ for all rows $i$.
2:For each row $i=1,\ldots,n$ in order:
  1. a.
    [1-CEx] For each column $j=1,\ldots,k$ in order: if $x_{ij}=+1$ or $n_{1,i}<c$, then change the sign of $x_{ij}$ if doing so improves the criterion.
  2. b.
    [2-CEx] Let $\mathcal{H}$ denote $\{j:x_{ij}=+1\}$ and let $\mathcal{L}$ denote $\{j:x_{ij}=-1\}$ using the current values in $X$. For each element $j\in\mathcal{H}$ in order, do the following:
    1. Find the column index $l\in\mathcal{L}$ for which swapping the signs of $x_{ij}$ and $x_{il}$ most improves the criterion. If such an improvement is possible, then swap the signs of $x_{ij}$ and $x_{il}$ and also replace $\mathcal{L}$ with ($(\mathcal{L}\cup\{j\})\setminus\{l\}$.
3:Repeat step 2 until no further exchanges can be made.
2.2 Computational Improvements
Weese et al. (2017) discuss how to quickly update the $E(s^{2})$ criterion when changing a coordinate in the design. Here we use a similar strategy while assuming $n_{1i}\leq c$. In steps 2a and 2b of the algorithm, we test if an exchange under consideration improves the criterion. If it does, then we update the design matrix, information matrix, and design criterion value accordingly. These updates are highly amenable to an implementation using vectorized operations, and our implementation also attempts to minimize the number of operations needed for each such test and update.
As before, we let $L=[\mathbf{1},X]$ and $S=L^{{}^{\prime}}L$. Defining $Q(X):={\mathop{\mathrm{tr}}}(L^{{}^{\prime}}LL^{{}^{\prime}}L)={\mathop{\mathrm{tr}}}(S^{2})$, we may write $UE(s^{2})=[Q(X)-n^{2}(k+1)]/(k(k+1))$. (Throughout this section and the next, we use the notation $A_{:,j}$ to represent the $j$th column of a matrix $A$.) We need to determine how much $Q(X)$ changes as the result of either a single-coordinate or two-coordinate exchange in $X$. For notational convenience in indexing, we use $x_{i0}=1$ to denote entries in the first column of $L$. This allows us to write
$Q(X)=\sum_{j=0}^{k}\sum_{l=0}^{k}s_{jl}^{2}=\sum_{j=0}^{k}\sum_{l=0}^{k}\bigg{(}\sum_{h=1}^{n}x_{hj}x_{hl}\bigg{)}^{2}$
Our test-and-update procedure is based on the following observation.
Lemma 1.
Consider two design matrices $X,\tilde{X}\in\{\pm 1\}^{n\times k}$ with $L=[\mathbf{1},X]$ and $S=L^{\prime}L$. If $X$ and $\tilde{X}$ differ only in row $i$, then for $J:=\{j=0,\ldots,k\mid\tilde{x}_{ij}=-x_{ij}\}$ we have
$Q(\tilde{X})-Q(X)=8|J|(k+1-|J|)-8\sum_{j\in J}\bigg{[}x_{ij}\sum_{l\not\in J}s_{jl}x_{il}\bigg{]}.$
The proof of this result is given in the Supplementary Materials. The lemma provides a simple test to determine if switching the signs of all $x_{ij}$ for fixed $i$ and $j\in J$ improves the criterion. The following result specializes it to the 1-coordinate and 2-coordinate exchanges in the proposed algorithm.
Corollary 1.
Consider a design matrix $X\in\{\pm 1\}^{n\times k}$ with $L=[\mathbf{1},X]$ and $S=L^{\prime}L$ and let $L_{i[j]}$ denote the row vector given by replacing element $j$ in the row vector $L_{i:}$ with a zero. (Recall that the columns of $L$ are indexed from 0.)
  1. a.
    Swapping signs on $x_{ij}$ decreases $Q(X)$ if and only if $k<x_{ij}L_{i[j]}S_{:,j}$. In this case, the change in $Q(X)$ will be $8(k-x_{ij}L_{i[j]}S_{:,j})$ and the corresponding rank-2 update of $S$ is given by subtracting $2x_{ij}L_{i[j]}$ from row $j$ of $S$ and subtracting its transpose from column $j$.
  2. b.
    Suppose $x_{ij}=+1$ and $x_{il}=-1$. Swapping signs on both $x_{ij}$ and $x_{il}$ decreases $Q(X)$ if and only if $2(k-1)+n<L_{i[j]}S_{:,j}-L_{i[j]}S_{:,l}+s_{jl}$. Among all such indices $l$, the improvement in the criterion will be greatest for those minimizing $L_{i[j]}S_{:,l}-s_{jl}$. In this case, the change in $Q(X)$ will be $8[2(k-1)+n-L_{i[j]}S_{:,j}+L_{i[j]}S_{:,l}-s_{jl}]$. We update $S$ by subtracting $2x_{ij}L_{i[j]}$ from row $j$ of $S$, adding $2x_{il}L_{i[l]}$ to row $l$, and also applying the corresponding column updates.
Notice that Corollary 1 allows us to test for decrease without updating $S$. It also shows that we can vectorize the $l$-index selection for 2-coordinate exchange by calculating the row vector $L_{i[j]}S_{:,\mathcal{L}}+S_{j,\mathcal{L}}$ and finding its minimum entry.
3 Properties of CRowS Designs
In this section, we study the designs we described in Section 2. In particular, we provide some initial bounds on the $UE(s^{2})$ criterion for the row-constrained designs, under the assumption that the row constraints are tight, which provides a design optimality check. Then, we empirically investigate several relevant design settings to better understand how the designs change as a function of the row constraint, providing insights into the practical effect of the constraints on the effectiveness of the row-constrained designs.
3.1 Some Bounds for CRowS Designs
In this section we adapt some ideas from Liu and Hickernell (2002). Recall the notation $Q(X):={\mathop{\mathrm{tr}}}(L^{{}^{\prime}}LL^{{}^{\prime}}L)$ and $L=[\mathbf{1},X]$ from §2.2. It is easily verified that $Q(X)=n^{2}+2\|X^{\prime}\mathbf{1}_{n}\|_{2}^{2}+{\mathop{\mathrm{tr}}}(X^{\prime}XX^{\prime}X)$.
Theorem 1.
We have
$Q(X)=n^{2}(1-k^{2})+2\|X^{\prime}\mathbf{1}_{n}\|_{2}^{2}+2n\|X\mathbf{1}_{k}\|_{2}^{2}+\sum_{l=1}^{k}\sum_{j=1}^{k}\|X_{:,l}-X_{:,j}\|_{1}^{2}.$ (7)
If $X\mathbf{1}_{k}=(2c-k)\mathbf{1}_{n}$ then the terms in (7) satisfy
$\|X\mathbf{1}_{k}\|_{2}^{2}=n(2c-k)^{2},$ (8)
$\|X^{\prime}\mathbf{1}_{n}\|^{2}_{2}\geq(k-\delta)(n-2\gamma)^{2}+\delta(n-2\gamma-2)^{2},$ (9)
$\sum_{l=1}^{k}\sum_{j=1}^{k}\|X_{:,l}-X_{:,j}\|_{1}^{2}\geq 4[(k^{2}-k)\phi^{2}+\psi(2\phi+1)]$ (10)
for integers $\gamma:=\lfloor nc/k\rfloor$, $\delta:=nc-k\gamma$, $\phi:=\lfloor 2nc(k-c)/(k^{2}-k)\rfloor$, and $\psi:=2nc(k-c)-(k^{2}-k)\phi$.
If the bound (9) is attained, then all the column sums for $X$ take values from $\{\gamma,\gamma+1\}$ in specific proportions. Likewise, if the bound (10) is attained, then all $l^{1}$-differences of columns in $X$ take values from $\{2\phi,2\phi+2\}$. That raises several questions. When must there exist feasible $X$ with all column sums in $\{\gamma,\gamma+1\}$? Must there exist feasible $X$ with all $\|X_{:,l}-X_{:,j}\|_{1}$ in $\{2\phi,2\phi+2\}$? Can both be satisfied simultaneously? We leave the study of these questions to future work.
3.2 Empirical Investigation of Row Constraint
An additional question, unanswered as of yet, is whether there is a threshold for $c$ below which we are guaranteed that a row-constrained $UE(s^{2})$-optimal design will have tight row constraints. In this section we investigate this question empirically. Reasoning intuitively, unconstrained $UE(s^{2})$-optimal designs try to make $s_{jl}$’s small, which means that the columns are pushed toward balance in order that the $s_{0j}$’s are small. In an unconstrained design, we therefore expect $+1$’s and $-1$’s should each occur approximately $nk/2$ times. Thus, for $c\geq k/2$, it is possible for the numbers of $+1$’s and $-1$’s to occur the same number of times, and we may observe slack in constraint (6) with the optimal $UE(s^{2})$ value similar to the unconstrained case. For $c<k/2$, however, we may see tight row constraints and, as $c$ gets smaller, an increasingly large difference between the row-constrained optimal $UE(s^{2})$ value and that of the unconstrained optimal design. That is, if $U^{*}$ is the $UE(s^{2})$ value for the unconstrained $UE(s^{2})$-optimal design and $U_{c}^{*}$ is the $UE(s^{2})$ value of the row-constrained $UE(s^{2})$-optimal design with $c$ $+1$’s per row, then $U^{*}\leq U_{c}^{*}$ if $c\leq k/2$, but as $c\rightarrow k$, $U_{c}^{*}\rightarrow U^{*}$.
To explore the relationship between $c$ and $UE(s^{2})$, along with the effect of $c$ on the tightness of the constraints in (6), we first examine a set of approximately $UE(s^{2})$-optimal designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, each constructed using $100$ algorithm tries of the algorithm described in Section 2.1. We consider several properties of the designs, each as a function of $c$: the $\text{UE}(s^{2})$ criterion value, along with the minimum, mean, and maximum slack in the row constraint, measured for each row as $(2c-k)-\sum_{j=1}^{k}x_{ij}$. We also plot the bounds from Section 3.1 for those designs without row constraint slack, and we provide some simulation results for various effect sizes. For the simulations, we generate $1{,}000$ datasets from model (1) with $\sigma^{2}=1$, $a=1$ nonzero element of $\bm{\beta}$ with value $\beta$, in line with application-specific sparsity levels of 1% or less. Since in these applications we know the desired effect directions, we consider $\beta\in\{0.375,0.5,0.75,1\}$, which translate to mean response differences (when the active compound is present versus absent) of $0.75$, $1$, $1.5$, and $2$, respectively. The datasets are analyzed using the Lasso as in (2). To be precise, we center $\mathbf{y}$ as $\mathbf{y}_{c}$ while also centering and scaling $X$ as $X_{cs}$ so that each column has mean 0 and the same length. The resulting Lasso estimates are denoted $\hat{\bm{\beta}}_{cs}$ with regularization parameter $\lambda_{cs}$. To reduce false positive rates, we perform a thresholding procedure in the following way. We assume known $\sigma^{2}$ (see discussion in Section 6 for a relaxation of this assumption) and, after obtaining $\hat{\bm{\beta}}$ for all $\lambda_{cs}$, we set any elements of $\hat{\bm{\beta}}_{cs}$ to 0 if they are less than $\sigma/8$, a value was chosen via testing to control FPR. For the resulting set of nonzero Lasso estimates, we obtain ordinary least squares estimates and choose the $\lambda_{cs}$ to minimize the Bayesian Information Criterion. The resulting set of nonzero estimates of $\bm{\beta}_{cs}$ constitute the set of factors deemed potential hits. The true positive rate (TPR) is the proportion of truly active factors that are correctly identified, and the false positive rate (FPR) is the proportion of truly inactive factors that are incorrectly deemed to be hits.
Figure 1 provides a visualization of $\sqrt{UE(s^{2})}$ and the measures of row-constraint slack for all $c$ (left) and for a smaller set of $c$ (right). In Figure 2, we provide simulation results as a function of $c$. We observe that between $c=60$ and $c=70$ the constraint ceases to measurably impact the criterion value, the amount of row-constraint slack, and the TPR. We also see that while the bounds appear to be fairly tight (Figure 1, top left), closer inspection (Figure 1, top right) shows that there is a substantial gap as $c$ approaches the value beyond which there will be row-constraint slack. Thus, we conclude that for these designs any row constraint larger than $k/2=72$ will not have an appreciable effect on the results of an experiment using them, though the constraint includes slack beginning at some value $c$ slightly smaller than $k/2$. We have seen this same pattern for a number of other design sets as well (see Supplementary Material). The simulation (Figure 2) also shows that even for $c$ as small as 5 or 10, the designs reliably categorize active and inactive factors for sufficiently large effect sizes.
Figure 1: Characteristics of CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$. Plots are a function of $c$, the number of $+1$’s allowed in each row of design. Top left: $UE(s^{2})$ value for each design (blue) along with the bounds from Section 3.1 (for all designs with no row-constraint slack). Bottom left: $(2c-k)-\sum_{j=1}^{k}x_{ij}$; the solid line is the average slack, while the lower and upper dashed lines are the minimum and maximum slack. Top right: The same as the top left plot, for $40\leq c\leq 84$. Bottom right: The same as the bottom left plot, for $40\leq c\leq 84$.
Figure 2: Simulation results for CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, for four effect sizes, assuming one active factor; the solid lines are true positive rates while the dashed lines are false positive rates.
4 Comparison of Biological Screening Methods
Most commonly, biological screening is conducted using a simplistic one-compound-one-well (OCOW) approach (Malo et al., 2006; Kainkaryam and Woolf, 2009; Malo et al., 2010). Alternatively, a non-statistical pooling strategy has been proposed (Kainkaryam and Woolf, 2008) (poolHiTS). In this section, we compare via simulation these existing methods with CRowS, our proposed procedure, which uses the designs of Section 2, along with the Lasso, to estimate the set of hit compounds. We have compared the approaches with regard to their statistical performance, measuring both the true positive rate (TPR, proportion of truly active factors identified as active) and false positive rate (FPR, proportion of truly inactive factors declared active).
4.1 Methods to Compare
For the sake of the comparative results in this section, we make several simplifying assumptions. First, we assume that we know the effect directions of interest. For instance, in our metallo-$\beta$-lactamase application, the goal is the inhibition of enzyme activity and so we look for a negative effect direction. Without loss of generality, we assume a positive effect direction in our simulations. Secondly, we assume a known ambient level of variability, $\sigma^{2}$, from model (1). This is a strong assumption but simplifies the analyses, particularly for OCOW and poolHiTS. Third, we assume a known background mean response, $\mu$, when there are no hits in a particular well. (We provide an alternative set of simulations in the Supplementary Material that relaxes the assumptions about $\mu$ and $\sigma^{2}$.) Finally, we assume that responses, or some standard transformation of them, can be reasonably approximated by a normal distribution.
4.1.1 One-Compound-One-Well (OCOW)
The first approach is the most straightforward: apply a single compound to a single well and evaluate whether the result is unexpected in the assumed effect direction. We are assuming normal responses, a known positive effect direction, and known values of $\sigma^{2}$ and $\mu$. So the analysis is quite simple: declare as active any compound whose response exceeds $\mu+F^{-1}(0.95)\sigma$, where $F^{-1}$ is the inverse normal distribution function.
4.1.2 poolHiTS
The poolHiTS procedure, using Shifted Transversal Designs (STD, Kainkaryam and Woolf, 2008), pools compounds as we suggest and also provides for row constraints. However, while its design and analysis approaches are intuitive, the STD is neither constructed using a statistically motivated criterion nor analyzed using a statistical model. In addition to this, they suffer from the following additional disadvantages compared to our proposed procedure. First, due to their cyclic construction, STDs are not available for many combinations of the numbers of compounds screened, wells and compounds per well. For example, to screen $k=96$ compounds with $n=96$ wells, one cannot construct an STD design with $c=30$ compounds per well. To obtain a higher $c$, one must decrease $n$. In our comparisons in Section 4, we had to choose $(n,k,c)$ values that STD allowed, rather than the most natural values. Real experimental screens will invariably have unique requirements, which limits the usefulness of poolHiTS. Second, the poolHiTS methodology assumes a binary response vector (that is, each well can be categorized as a “hit” if it contains an active compound, or a “miss” if it doesn’t), so in the standard numeric-response setting it is required to determine a threshold beyond which a well’s response is categorized as a “hit”.
Details can be found in Kainkaryam and Woolf (2008), but here we briefly describe the basic poolHiTS methodology. The user specifies the total number of compounds $k$ and the maximum number of compounds per well $c$, along with the anticipated maximum number of active compounds and an overall error rate. This allows the construction of a design with $a$ layers, each layer consisting of $q$ pools, where the parameters $a$ and $q$ are determined by the procedure. This results in a final STD with $n=aq$ wells, which highlights its lack of flexibility to choose the number of wells for a given $k$ and $c$. The decoding algorithm requires a binary response vector; thus, to account for uncertainty associated with such a binarization, we specify a threshold, much like that for OCOW, which we can use to label each well. We conducted some tests to establish a threshold which would best balance TPR and FPR. In the end, assuming a positive effect direction, we chose $\mu+F^{-1}(0.96)\sigma$, where $F^{-1}$ is again the inverse standard normal distribution function; that is, any wells whose response fell beyond the threshold were labeled “hits” and others assumed inert. Once the pools have been labeled, the decoding algorithm proceeds by eliminating compounds if they appear in at least $E+1$ inert wells, and tagging compounds as active if they aren’t labeled as inert and appear in at least $E+1$ positive pools. Here $E$ is the number of errors allowed in the analysis. We note that it is possible for some compounds to fall into a third inconclusive category, and in this case we label these as active since further testing would be required to eliminate them.
4.1.3 Constrained Row Screening
Our proposed approach uses a CRowS design from Section 2, constructed using 100 algorithm randomized starts of the procedure described in Section 2.1, and then analyzed using the Lasso. For the simulations, where using a profile plot to choose active effects is impractical, we use a thresholding procedure to reduce FPR, with the threshold chosen via informal testing. In particular, our analysis procedure is:
  • Center and scale the design $X$ so that each column has the same length, call it $X_{cs}$; and center response $\mathbf{y}$, call it $\mathbf{y}_{c}$;
  • Obtain Lasso solutions for $\log(\lambda)$ ranging from $-8$ to $\log(\max(|X_{cs}^{T}\mathbf{y}_{c}|))$, the range chosen to reasonably ensure a full exploration of the Lasso estimation profile.
  • Across all $\lambda$, set any parameter estimate to 0 if it is smaller than $\sigma/8$. Because we assume a known effect direction, we threshold all negative estimates.
  • Refit each model, across $\lambda$’s, using least squares, and choose the model with the smallest BIC. The factors with nonzero estimates for this model are declared hits.
4.2 Description of Simulation Protocol
For the two pooling procedures (STD and SSD), data are simulated from the model in Eq. (1), assuming without loss of generality that $\sigma^{2}=1$. For the designs $X$ in the model, we used a set of nine different designs, with $(n,k,c)$ combinations as given in Table 1. To more closely mimic conditions in a lab setting, we preferred to use a different set of designs—all nine combinations of ($n=96,k\in\{96,150,192\},c\in\{10,30,50\}$)—but since STDs of these preferred sizes cannot be constructed, we accommodated the poolHiTS procedure by using the designs in Table 1. For the effect sizes $\bm{\beta}$, we considered effect sizes, in units of $\sigma$, of $D\in\{0.75,1,1.5,2,2.25,3,4\}$; that is, $D$ gives the difference in mean response between a pool which includes a hit compound and one that does not. Since our designs are parameterized such that $-1$ and $1$ represent the two levels, the elements of $\bm{\beta}$ are either 00 or $D/2$. Also the parameterization leads to $\beta_{0}=\mu+a\times\frac{D}{2}$, where $\mu$ is the background response mean and $a$ and the number of active factors. We fix $a=1$ for all simulations in this section, as this represents between $0.5\%$ and $1\%$ of total compounds studied.
Table 1: Design sizes for simulations. Both STD and CRowS designs were constructed for each of the sizes.
n k c n k c n k c
88 96 10 85 150 10 92 192 10
88 96 30 91 150 30 99 192 30
88 96 50 91 150 50 99 192 50
4.3 Comparison of CRowS with Alternatives
Here we provide a comparison of the three methods of Section 4.1, labeled OCOW, poolHiTS, and CRowS, in Figure 3. Since we are assuming known $\sigma^{2}$, the OCOW results will be unchanged across the nine design scenarios we consider up to variation across simulations, so we have only displayed it for the most difficult setting, ($n=96,k=192,c=10$).
In Figure 3(a), we clearly see the advantage of CRowS in terms of its ability to identify active compounds, compared to OCOW and poolHiTS. We also notice that the most restrictive constraint setting ($c=10$) results in a clear reduction in TPR, compared to less restrictive $c$ values. Furthermore, Figure 3(b) shows that poolHiTS fails to control FPR for most of the settings simulated. In highly sparse systems where thousands of compounds are studied, large FPRs could mean dozens or even hundreds of extra follow-up experiments on ultimately inert factors.
(a) True Positive Rates for the three methods.
(b) False Positive Rates for the three methods.
Figure 3: Comparisons of three design/analysis approaches to biological screening: One-Compound-One-Well (OCOW), the proposed method (CRowS), and the poolHiTS procedure with shifted transversal designs (poolHiTS). For CRowS and poolHiTS, the number of wells are specified in Table 1; for OCOW, the number of wells is equal to $k$.
5 Real Experiments
Metallo-$\beta$-lactamases exist in bacteria that cause nosocomial infections, including secondary bacterial infections contributing to an estimated 50% of COVID-19 deaths (Mojica et al., 2022; Chedid et al., 2021). Thus, the search for inhibitors of these enzymes is of considerable interest. To provide proof-of-concept of the CRowS approach, we performed four physical experiments for the inhibition of Verona Integron-encoded Metallo-$\beta$-lactamase 2 (VIM-2). Each of these experiments consisted of 1 known inhibitor and 30 known inert compounds, under a variety of conditions (Figure 4). The first experiment included no row constraints and an n/k ratio near 1. The other three experiments were more difficult in various ways, whether a smaller n/k ratio (Experiment 2) or row constraints (Experiments 3 and 4). The designs for each experiment were constructed using the procedure described in Section 2; they are available in the Supplementary Material. The compounds were arrayed into normalized pools implied by the designs such that each compound was at a concentration of 30 $\mu$M. We transferred the compounds into the pools using an Eppendorf epMotion 5073m liquid handing workstation. Enzymatic assays were carried out in microplates at 25C and results were determined by reading the absorbance at 495 nm in a Molecular Devices SpectraMax iD5 plate reader. The list of compounds, the four designs, and the experimental results are provided in the Supplementary Material.
Figure 4 shows the Lasso estimate profile plots for the four experiments. In the first three experiments, L-captopril is clearly identified as an inhibitor, though in Experiments 1 and 3 they also identify a possible false positive, Adenosine-5’-triphosphate disodium salt hydrate (abbreviated A-5’-dsh). The identification of 1 false positive out of 30 inert compounds is in keeping with a false positive rate of around 3%. Note that in Experiment 2 we do not consider 1,6-Hexanediol (abbreviated 1,6-Hxnd) as a false positive because its potential effect is not inhibitory. We see that even in experiment 3, in which the design limits the number of compounds per well to 10, we clearly identify the known inhibitor. On the other hand, in the fourth experiment, where only five compounds per well are allowed, we fail to identify the known inhibitor though we also identify no false positives.
Figure 4: Profile plots for four proof-of-concept experiments. For each plot, the largest five effects are annotated. The experiments are identified by $(n,k,c)$.
6 Discussion
In this paper, we have developed a new class of supersaturated designs, motivated by applications in high-throughput screening which use multiwell plates to search for promising chemical compounds. These designs constrain the number of compounds included in each well based on logistical and biological concerns of the specific HTS setting. These designs are used as part of a method we call CRowS that use both a statistical criterion—the $UE(s^{2})$-criterion—to guide construction of the design and a statistical analysis method—the Lasso—to obtain the set of possible hit compounds. Thus, our work provides a novel and improved solution to a scientific problem as well as innovative statistical design methodology inspired by the problem. In addition, we have demonstrated an important application of supersaturated designs, including real experimental results that display the proposed methodology in action. Finally, though the proof-of-concept experiments were small, we show via simulation the use of larger SSDs than have previously been typically studied, and our methodology can scale to the hundreds or thousands of factors, opening up a rich set of computational and statistical problems for the statistical community to explore.
More specifically, in this work we have developed an efficient computational procedure to construct CRowS designs, which minimize the $UE(s^{2})$ criterion subject to $c$, the number of compounds allowed per well. We have also studied the behavior of these designs as a function of $c$, both theoretically and empirically. Theoretically, we provide initial lower bounds on the $UE(s^{2})$ criterion which can be used in some cases to verify that the globally optimal design has been obtained. Empirically, we find that when $c$ is between 40% and 45% of $k$, the constrained design is not appreciably worse than the unconstrained design. Furthermore, even severely constrained designs with $c$ as small as 5 or 10 still reliably find hit compounds in simulations. We also find that compared to the standard one-compound-one-well practice, or a non-statistical method that allows the pooling of multiple compounds in a well, CRowS detects hit compounds with substantially more reliability, while controlling the rate of false discoveries.
Under the settings considered, our approach represents a clear improvement over existing HTS design and analysis methodology. However, there are several limitations to our work as well. First, though our design construction procedure is fast for the sizes of designs reported in this paper (on the order of seconds) and also for designs with $n$ and $k$ both in the hundreds (on the order of minutes), when $n$ and $k$ are both in the thousands the procedure takes on the order of days to produce a locally optimal design. This may not be a major concern because even a screening campaign studying thousands of compounds could be broken down into designs built on 96- or 384-well plates. If, however, screeners wished to experiment on, say, 1,536-well plates, they would need to be willing to wait several days for a design. Related to this are liquid handling challenges. One advantage of one-compound-one-well is that only a limited amount of liquid must be transferred from source to target plates. If instead of a single compound per well, we have $c$ compounds per well, then the time it takes to prepare the plates will increase proportionally to $c$. This, of course, must be weighed against the improved hit rates that pooling using our method could provide.
Another important concern is interactions. In this work, we have assumed that compounds don’t interact either statistically or chemically. A natural question is whether statistical interactions between factors might affect the CRowS method’s ability to detect hits. In order to investigate this we repeated the simulations described in Section 4 where the response was generated using the model in (1) but the $\bm{\beta}$ included two active factors and a single active interaction. We considered both synergistic and antagonistic interactions that obey strong (interaction only appears when both parents are active) or weak (interaction only includes one active parent) heredity. In the Supplementary Material, we provide results of this simulation. We find that synergistic interactions degrade the ability of CRowS to detect hit compounds, while antagonistic interactions improve the results. We do note that Singh and Stufken (2023) has recently proposed a method to estimate interactions in SSD’s. We are currently working to adapt and apply Singh and Stufken (2023) to the present setting, so that important interactions can be accounted for and/or estimated.
On the application side, our work leads to a number of interesting questions. Though there are numerous examples of successful pooling in the literature (e.g. Devlin et al., 1996; Wilson-Lingardo et al., 1996; Taylor et al., 1997) and even in the context of commercial drug discovery (Elkin et al., 2015), there is also a concern regarding so-called “promiscuous interactions” (Feng and Shoichet, 2006), in which compounds may react to each other in unpredictable ways. In our own experimental work, before fine-tuning the process by which the designs were implemented we twice failed to achieve successful results. In the first attempt in particular, we were experimenting on a wider set of compounds than the group we have reported on in this paper. Additional experiments may clarify from the chemical side the limits of pooling in this or other contexts.
This work raises a number of theoretical questions as well. From our empirical work, it appears that row-constraint slack begins to appear at between 40% and 45% of $k$. Can this be characterized more precisely? Can we specify a threshold for $c$ beyond which we are guaranteed that we will have no row-constraint slack? Can we tighten the bounds on the designs, and perhaps develop construction methods for certain $(n,k,c)$ combinations that bypass the need for an algorithm altogether?
Supplementary Material
  • A
    Supplementary document which includes proofs, additional results, and details regarding the real experiments.
  • B
    Matlab code to construct CRowS designs.
  • C
    Designs from the paper, excepting the designs for the real experiments.
  • D
    Designs, data, and analysis of the real experiments.
Acknowledgements, Funding, and Disclosure Statement
This work was funded in part by Miami University’s Faculty Research Committee (BJS) and by the National Institutes of Health through grant GM128595 (RCP). We also thank Jens Mueller from Miami’s Research Computing Support for his computational assistance. The authors report there are no competing interests to declare.

### v2

@@
 Abstract
-High-throughput screening, in which multiwell plates are used to test large numbers of compounds against specific targets, is widely used across many areas of the biological sciences and most prominently in drug discovery.
-We propose a statistically principled approach to these screening experiments, using the machinery of supersaturated designs and the Lasso.
-To accommodate limitations on the number of biological entities that can be applied to a single microplate well, we present a new class of row-constrained supersaturated designs.
-We develop a computational procedure to construct these designs, provide some initial lower bounds on the average squared off-diagonal values of their main-effects information matrix, and study the impact of the constraint on design quality.
-We also show via simulation that the proposed constrained row screening method is statistically superior to existing methods and demonstrate the use of the new methodology on a real drug-discovery system.
-Keywords: experimental design; biological screening; regularization
+High-throughput screening, in which large numbers of compounds are traditionally studied one-at-a-time in multiwell plates against specific targets, is widely used across many areas of the biological sciences including drug discovery.
+To improve the efficiency of these screens, we propose a new class of supersaturated designs that guide the construction of pools of compounds in each well.
+Because the size of the pools are typically limited by the particular application, the new designs accommodate this constraint and are part of a larger procedure that we call Constrained Row Screening or CRowS.
+We develop an efficient computational procedure to construct the CRowS designs, provide some initial lower bounds on the average squared off-diagonal values of their main-effects information matrix, and study the impact of the constraint on design quality.
+We also show via simulation that CRowS is statistically superior to the traditional one-compound-one-well approach as well as an existing pooling method, and demonstrate the use of the new methodology on a Verona Integron-encoded Metallo-$\beta$-lactamase 2 assay.
+Keywords: experimental design; biological screening; the Lasso
 1 Introduction
 High-throughput screening (HTS) of chemical or biological compounds is used extensively across a range of critical scientific enterprises.
-This type of screening is common in basic functional genomics studies and synthetic biology applications (Zhou et al., 2014; Zeng et al., 2020; Sarnaik et al., 2020), but arguably its highest impact application is drug discovery (Perola, 2010; Blay et al., 2020; Soheilmoghaddam et al., 2021; Aldewachi et al., 2021).
-The literature suggests that nearly one-third of clinical drug candidates emerge from high-throughput screening campaigns (Brown and Boström, 2018).
-These screens were a centerpiece in the National Cancer Institute’s Cancer Moonshot program (Wilson et al., 2020).
-But as liquid handling, miniaturization, and other screening technologies have improved, the number of compounds to explore has increased as well (Elkin et al., 2015; Volochnyuk et al., 2019).
-This, in turn, highlights the need to improve the experimental and statistical aspects of these screens.
-In the current work, we present a scalable statistical methodology, including a novel experimental design approach, to make high-throughput screening more flexible, cost-effective, and statistically efficient for detecting promising compounds, compared to existing methods.
-We do this by proposing a new type of supersaturated design which assigns a number of compounds to each well.
-This design is then analyzed using regularized regression.
-These new designs are characterized by a constraint which governs the number of compounds which can be included in each well.
-Using the unconditional $E(s^{2})$ supersaturated design criterion, we present a computationally efficient heuristic to construct these designs, and study them by developing lower bounds on the criterion and investigating how design quality changes as a function of the constraint.
-We also give evidence of the effectiveness of our approach via simulation.
-Finally, we demonstrate the methodology in real experimental screens used to identify an inhibitor of a metallo-$\beta$-lactamase.
+This type of experiment is common in basic functional genomics studies and synthetic biology applications (Zeng et al., 2020; Sarnaik et al., 2020), but arguably its highest impact application is drug discovery (Blay et al., 2020; Soheilmoghaddam et al., 2021; Aldewachi et al., 2021).
+The literature suggests that nearly one-third of clinical drug candidates emerge from high-throughput screening campaigns (Brown and Boström, 2018), and these screens were a centerpiece in the National Cancer Institute’s Cancer Moonshot program (Wilson et al., 2020).
+For example, HTS is used in the search for inhibitors of enzymes that cause antibiotic resistance.
 The prevalence of such enzymes is of great concern because they exist in bacteria that cause nosocomial infections, including in secondary bacterial infections contributing to an estimated 50% of COVID-19 deaths (Mojica et al., 2022; Chedid et al., 2021).
-Searching for compounds that inhibit these enzymes’ activity is an important part of the arsenal for addressing antibiotic resistance (CDC, 2019).
-HTS utilizes plates with varying well capacities (e.g., 96, 384, or 1536).
-Each well contains cells or enzymes and is subjected to application-specific conditions.
-Most commonly, a single chemical compound is assigned to a well, a biological or chemical process then occurs, and a measurement is taken.
-By using many plates, huge compound libraries can be studied.
-Compounds that produce meaningful measurements are then validated by more focused experimentation.
-The process is called screening because the goal is to detect, among thousands of candidates, the few compounds that exhibit the most promise.
-Though experimental design and statistical analysis have been promoted in HTS, and sophisticated methods of normalization and bias-correction have been developed (Altman, 2005; Birmingham et al., 2009; Malo et al., 2010; Dragiev et al., 2011; Zhang, 2011; Caraus et al., 2015), the basic experimental design approach most often used is simplistic: apply each treatment to one well (Malo et al., 2006; Kainkaryam and Woolf, 2009) (or, as recommended in the literature, in triplicate; see Malo et al., 2010) and then use a z-score to determine if the result is a hit.
-As shown in Section 4, this approach’s advantage of simplicity comes at the cost of reduced ability to detect hits.
-We note the recent work using machine/deep learning to preselect the “best” compounds (e.g., Stokes et al., 2020; Cheng et al., 2024), and these approaches have become popular for drug discovery (Mayr et al., 2018; Camacho et al., 2018).
-Our approach is complementary to these efforts, in the sense that it might be used to create training sets for HTS machine/deep learning models.
-We are not the first to propose pooling multiple compounds in each well; in fact, versions of pooling have been used successfully (see Kainkaryam and Woolf, 2009).
-The most aggressive version of the existing strategies is called non-adaptive pooling (Thierry-Mieg, 2006; Thierry-Mieg and Bailly, 2008; Wilson et al., 2020; Kainkaryam and Woolf, 2009), which purportedly can screen, with a high degree of accuracy, large numbers of compounds with relatively few pools.
-The most relevant approach uses so-called Shifted Transversal Designs (STDs) within a larger procedure called poolHiTS.
-This method requires that the response for each well can be identified as either a “hit” or “inert”, and also that the user specifies the expected number of hits and the expected error rate.
-STDs construct pools to have as little overlap between them as possible, and they use information about which pools trigger hits to back out which compounds are likely hits and which are inert.
-The chief advantage of STDs is that they are fast to construct, but they have several important drawbacks described in more detail in Section 4.1.2.
-Most basically, poolHiTS treats neither the design nor the analysis in a statistical way.
-Our approach, on the other hand, frames the problem as a statistically designed experiment, which generates data posited to arise from a statistical model.
-This allows us to use statistical estimation procedures and thereby control false positive rates with high true positive rates.
-Overall, we will show that our procedure is more flexible with better statistical performance than poolHiTS.
-Our approach uses a supersaturated design (SSD) to guide the placement of multiple compounds in each well, where $k$ compounds are investigated using $n$ wells (with $n\leq k$) while controlling false positive rates and increasing true positive rates.
-SSDs—experiments for which there are more factors than runs—were first described over sixty years ago (Satterthwaite, 1959; Booth and Cox, 1962).
-Motivated by the desire to experiment on complex high-cost systems, SSDs have seen extensive development over the last thirty years (e.g., Lin, 1993; Wu, 1993; Nguyen, 1996; Cheng, 1997; Cheng and Tang, 2001; Butler et al., 2001; Bulutoglu and Cheng, 2004; Georgiou, 2014; Jones and Majumdar, 2014; Weese et al., 2021).
-In our setting, an SSD is a two-level design composed of $k$ compounds (factors; absence with level $-1$; presence with level $+1$), and $n$ wells (runs) with $k\geq n$.
-The SSD provides a set of instructions regarding which compounds are included in which well.
-The design is then analyzed assuming the main-effects statistical model
-$\mathbf{y}=\beta_{0}\mathbf{1}+X\bm{\beta}+\bm{\epsilon},$ (1)
-where $X$ is the $n\times k$ SSD design matrix with elements $x_{ij}=\pm 1$, $\bm{\beta}=(\beta_{1},\dots,\beta_{k})^{T}$ and $\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}I)$.
-Because $k\geq n$, ordinary least squares cannot be used to estimate factor effects.
-Although forward selection approaches (Westfall et al., 1998; Abraham et al., 1999; Kelly III and Voelkel, 2000) were initially considered, regularization methods are now preferred for the analysis of SSDs.
-The Gauss-Dantzig selector (Candes and Tao, 2007) has been shown to be effective (Phoa et al., 2009; Marley and Woods, 2010; Weese et al., 2015), but the Lasso (Tibshirani, 1996) is closely related to the Dantzig selector (Meinshausen et al., 2007; Lounici, 2008; James et al., 2009; Asif and Romberg, 2010) and achieves similar performance (see Draguljić et al., 2014).
-Additionally, Stallrich et al.
-(2023) has recently developed an optimal screening design framework based on the Lasso.
-Thus, in our work we use the Lasso, which selects estimates according to
-$(\hat{\beta}_{0},\hat{\bm{\beta}})=\operatorname*{arg\,min}_{\beta_{0},\bm{\beta}}\frac{1}{2n}||\bm{y}-\beta_{0}\bm{1}-X\bm{\beta}||_{2}^{2}+\lambda||\bm{\beta}||_{1},$ (2)
-where $||\bm{\beta}||_{\ell}=(\sum_{j=0}^{k}|\beta_{j}|^{\ell})^{1/\ell}$ and $\lambda>0$.
-A critical aspect of the proposed methodology is constructing SSDs to use as maps guiding the assignment of compounds to wells.
-SSDs are classically constructed using the $E(s^{2})$ criterion (Booth and Cox, 1962; Nguyen, 1996; Cheng, 1997; Bulutoglu and Cheng, 2004) to create designs with columns balanced between $-1$’s and $+1$’s, as close to mutually orthogonal as possible.
-We follow work that relaxes the balance requirement (Marley and Woods, 2010; Jones and Majumdar, 2014; Weese et al., 2015), leading to the unconditional $E(s^{2})$ criterion.
+As liquid handling, miniaturization, and other screening technologies have improved, the number of potentially inhibitive compounds to explore has increased as well (Elkin et al., 2015; Volochnyuk et al., 2019), which motivates the need to improve the experimental and statistical aspects of these screens.
+Classically, a single compound is applied to a single well in a 96-, 384-, or 1536-well plate (Malo et al., 2006; Kainkaryam and Woolf, 2009), an assay is performed with relevant normalization and bias-correction (Altman, 2005; Birmingham et al., 2009; Malo et al., 2010; Dragiev et al., 2011; Caraus et al., 2015), and hit compounds are identified with a z-score.
+As an alternative to one-compound-one-well (OCOW), researchers have suggested creating pools of compounds in each well which can improve HTS because compounds are observed in multiple wells rather than just one.
+Various versions of pooling have been used successfully (e.g., Devlin et al., 1996; Wilson-Lingardo et al., 1996; motlekar2008evaluation; Kainkaryam and Woolf, 2009; Elkin et al., 2015; silva2016output), and the most aggressive version, called non-adaptive pooling (Thierry-Mieg, 2006; Kainkaryam and Woolf, 2009), is said to accurately screen large numbers of compounds with relatively few pools.
+However, these methods are largely non-statistical in both their design and analysis.
+In this work, we develop a novel class of row-constrained supersaturated designs that provides the pooling instructions to run a $k$-compound screen in $n$ wells, while accounting for a screener-specified constraint on the maximum number of compounds per well.
+The constraint may be due, for instance, to the size of the well, a required minimum compound concentration, or the maximum ionic strength of the solution beyond which the compounds may precipitate.
+We call these limitations row constraints, denoted by $c$, because they limit the number of $+1$’s in each design row.
+Designs are constructed using the unconditional $E(s^{2})$ criterion from the supersaturated design literature, and we analyze them using the Lasso.
+These designs are part of a larger procedure, which we call Constrained Row Screening, or CRowS, that includes both the design and its analysis via the Lasso.
+We demonstrate that altogether CRowS is more flexible, cost-effective, and statistically efficient than traditional methods.
+We present a computationally efficient heuristic to construct the designs of any relevant $(n,k,c)$ combination, study the designs both theoretically and empirically as a function of $c$, provide evidence that our methodology improves upon both the OCOW approach as well as the pooling method of Kainkaryam and Woolf (2008), and show in a set of real experiments that it can be used to identify an inhibitor of a particular antibiotic-resistant enzyme.
+Thus, our contributions are a statistically principled and improved solution to an important biological problem and innovation in the construction of the relevant designs.
+(We note that after our work was completed we became aware of a recent paper by ji2023target that constructs pools by optimizing a “sensing matrix” which also uses a version of the Lasso to identify hits.
+They don’t explicitly constrain the size of the pools, nor do they engage with the statistical literature or study their method theoretically or via simulation.
+Still, it is encouraging that the biomedical community is considering statistical approaches to solve high-throughput screening problems.)
+Our work emerges from research on supersaturated design (SSD), where $k$ factors are investigated using $n$ runs (with $n\leq k$).
+SSDs, which were first described over sixty years ago (Satterthwaite, 1959; Booth and Cox, 1962), have seen extensive development over the last thirty years (e.g., Lin, 1993; Wu, 1993; Nguyen, 1996; Cheng, 1997; Georgiou, 2014; Jones and Majumdar, 2014).
+In our setting, an SSD is a two-level, $n\times k$ design $X$, composed of $k$ compounds (factors; absence with level $-1$; presence with level $+1$), and $n$ wells (runs) with $k\geq n$ (along with a constraint on the design rows, described below and in Section 2).
+The SSD provides a set of instructions regarding which compounds are included in which well, so one may improve the design by constructing it using a statistically relevant criterion.
+SSDs are classically constructed using the $E(s^{2})$ criterion (Booth and Cox, 1962; Nguyen, 1996) to create balanced columns, as close to mutually orthogonal as possible.
+The row constraint of Section 2 typically necessitates a design with more $-1$’s than $+1$’s, so we use the unconditional $E(s^{2})$ criterion (Marley and Woods, 2010; Jones and Majumdar, 2014; Weese et al., 2015) which has the same orthogonality goal without the requirement of column balance.
 Mathematically, let $L=[\mathbf{1},X]$, and $S=L^{{}^{\prime}}L$ with elements $s_{ij}$ and $O_{S}=\sum_{i<j}s_{ij}^{2}$.
@@
 the average of the squared off-diagonal elements of $S$.
-In many cases $UE(s^{2})$-optimal designs can be constructed directly from known Hadamard matrices (Jones and Majumdar, 2014), but in our work we construct the designs algorithmically; see Section 2.1.
-Our approach, which we call constrained row screening, or CRowS, breaks new ground in several important ways.
-First, it provides a statistically principled and substantially improved solution to an important problem in biology.
-We will demonstrate empirically in Section 4 that our approach has much better statistical properties, in terms of true and false positive rates, than existing methods.
-Secondly, the constraints inherent to the HTS application necessitate innovation in design construction.
-Specifically, though there is literature suggesting that multiple compounds can be applied to individual wells (e.g., Elkin et al., 2015; Kainkaryam and Woolf, 2008, 2009), it is also clear that in drug screening settings, and by extension high-throughput screening more broadly, there are limitations regarding the number of compounds that can be added to each well.
-These may be due to the size of the well, the minimum concentration required in order for a compound to be detectable, or the maximum ionic strength of the solution beyond which the compounds may precipitate.
-We call these limitations the row constraints, denoted by $c$, because it limits the number of $+1$’s in each row of the design.
-Thus, we have developed a new coordinate-exchange algorithm (Meyer and Nachtsheim, 1995) that aims to minimize $UE(s^{2})$ while respecting row constraints, and we study these new CRowS designs both theoretically and empirically as a function of $c$.
-Third, our work provides a significant, broad-ranging application for SSDs, which has been lacking in the literature.
-Despite their considerable promise (Weese et al., 2021), we only know of a handful of scattered articles that report on their use (Carpinteiro et al., 2004; Dejaegher and Vander Heyden, 2008; Jridi et al., 2015; Zarkadas and Besseris, 2023).
-Our work changes that, by applying these designs to a large class of biological/chemical screens and demonstrating their usefulness in real experiments.
-Finally, SSD research has focused largely on designs with at most dozens of factors; in contrast, CRowS designs will necessarily be larger in order to apply them to HTS.
-The designs we investigate in this paper have more than one hundred factors, though the proof-of-concept real experiments we report on in Section 5 are smaller.
-Since HTS studies routinely run into the tens of thousands and even much larger (Elkin et al., 2015), the size of the SSDs applied to this area will be limited not by the number of compounds studied, but by, in large part, logistical constraints related to the liquid handling automation (see Section 5 for a discussion) as well as computational limitations in the construction of large designs.
-We note that Eskridge et al.
-(2004) study large SSDs, and even given a nod to HTS as an application, but they don’t provide designs for $n>24$ nor do they consider row constraints or provide a real example.
+In many cases $UE(s^{2})$-optimal designs can be constructed directly (Jones and Majumdar, 2014), but for our row-constrained versions, we construct the designs algorithmically; see Section 2.1.
+After the design is constructed and executed, it is analyzed assuming the main-effects model
+$\mathbf{y}=\beta_{0}\mathbf{1}+X\bm{\beta}+\bm{\epsilon},$ (2)
+where $\bm{\beta}=(\beta_{1},\dots,\beta_{k})^{T}$ and $\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}I)$.
+Because $k\geq n$, ordinary least squares cannot be used to estimate factor effects and regularization methods are preferred for analysis.
+The Gauss-Dantzig selector (candes_tao2007) is effective (Phoa et al., 2009; Marley and Woods, 2010; Weese et al., 2015), while the closely related Lasso (Tibshirani, 1996; Meinshausen et al., 2007; Lounici, 2008; James et al., 2009; Asif and Romberg, 2010) achieves similar performance (see Draguljić et al., 2014).
+Stallrich et al.
+(2023) has also recently developed an optimal screening design framework based on the Lasso.
+Thus, to analyze our proposed designs we use the Lasso, which selects estimates according to
+$(\hat{\beta}_{0},\hat{\bm{\beta}})=\operatorname*{arg\,min}_{\beta_{0},\bm{\beta}}\frac{1}{2n}||\bm{y}-\beta_{0}\bm{1}-X\bm{\beta}||_{2}^{2}+\lambda||\bm{\beta}||_{1},$ (3)
+where $||\bm{\beta}||_{\ell}=(\sum_{j=0}^{k}|\beta_{j}|^{\ell})^{1/\ell}$ and $\lambda>0$.
 The remainder of the paper is structured as follows.
-In Section 2, we describe row-constrained supersaturated designs in more detail, along with the algorithm we use to construct them.
-Section 3 provides some initial lower bounds on the $UE(s^{2})$ values for these CRowS designs, as well as an empirical study that sheds light on their behavior as a function of the severity of the row constraint.
-In Section 4 we use a set of extensive simulations to show the effectiveness of our proposed procedure compared to alternative methods.
+In Section 2, we describe CRowS designs in more detail, along with the algorithm we use to construct them.
+Section 3 provides initial lower bounds on the $UE(s^{2})$ values for these CRowS designs, as well as an empirical study that sheds light on their behavior as a function of the severity of the row constraint.
+In Section 4 we use a set of simulations to show the effectiveness of our proposed procedure compared to alternative methods.
 In Section 5 we provide more detail on the metallo-$\beta$-lactamase system, and describe the results of several screens as a real-world proof-of-concept of CRowS.
@@
 2 Row-Constrained Supersaturated Designs
-In this section we describe the new class of supersaturated designs that have emerged from this work, along with the heuristic optimization procedure we have developed to construct them.
+In this section we describe the new class of supersaturated designs that we propose, along with the heuristic optimization procedure developed to construct them.
 As explained in Section 1, $UE(s^{2})$-optimal designs minimize the average of the squared off-diagonal elements of information matrix $S$.
 In the present setting, we propose a generalization of $UE(s^{2})$-optimal designs that impose a constraint on the number of $+1$’s in each row.
-Thus, using notation from Section 1, we say that an $n\times k$ design $X$ is row-constrained $UE(s^{2})$-optimal if $n\leq k$, $[X]_{ij}=x_{ij}$, and the design solves the following optimization problem:
+Thus, using notation from Section 1, we say that an $n\times k$ design $X$ is a CRowS design if $n\leq k$, $[X]_{ij}=x_{ij}$, and the design solves the following optimization problem:
 min $UE(s^{2})$ (4)
@@
 where $UE(s^{2})$ is defined in (1) and $c$ is the maximum number of compounds/well.
-As mentioned in the Introduction, we call the proposed row-constrained $UE(s^{2})$-optimal designs constrained row screening, or CRowS, designs.
 2.1 Construction Algorithm
@@
 Let $n_{1i}$ be the number of $+1$’s in design row $x_{i}$, so that $n_{1i}\leq c\;\forall i$.
-The algorithm is described in Algorithm 1.
+The algorithm construction is described in Algorithm 1.
 For the designs in this paper, we ran Algorithm 1 with 100 random starts and chose the design with the best objective function value.
@@
 Our test-and-update procedure is based on the following observation.
-Lemma 1.
+Proposition 1.
 Consider two design matrices $X,\tilde{X}\in\{\pm 1\}^{n\times k}$ with $L=[\mathbf{1},X]$ and $S=L^{\prime}L$.
@@
 $Q(\tilde{X})-Q(X)=8|J|(k+1-|J|)-8\sum_{j\in J}\bigg{[}x_{ij}\sum_{l\not\in J}s_{jl}x_{il}\bigg{]}.$
-The proof of this result is given in the Supplementary Materials.
-The lemma provides a simple test to determine if switching the signs of all $x_{ij}$ for fixed $i$ and $j\in J$ improves the criterion.
+The proof of this result is given in Section 1 of Supplementary Material A.
+The proposition provides a simple test to determine if switching the signs of all $x_{ij}$ for fixed $i$ and $j\in J$ improves the criterion.
 The following result specializes it to the 1-coordinate and 2-coordinate exchanges in the proposed algorithm.
@@
 3 Properties of CRowS Designs
-In this section, we study the designs we described in Section 2.
+In this section, we study the designs described in Section 2.
 In particular, we provide some initial bounds on the $UE(s^{2})$ criterion for the row-constrained designs, under the assumption that the row constraints are tight, which provides a design optimality check.
@@
 3.1 Some Bounds for CRowS Designs
-In this section we adapt some ideas from Liu and Hickernell (2002).
+The result in this section uses ideas adapted from Liu and Hickernell (2002); it is proven in Section 2 of Supplementary Material A.
 Recall the notation $Q(X):={\mathop{\mathrm{tr}}}(L^{{}^{\prime}}LL^{{}^{\prime}}L)$ and $L=[\mathbf{1},X]$ from §2.2.
@@
 3.2 Empirical Investigation of Row Constraint
-An additional question, unanswered as of yet, is whether there is a threshold for $c$ below which we are guaranteed that a row-constrained $UE(s^{2})$-optimal design will have tight row constraints.
+An additional unanswered question is whether there is a threshold for $c$ below which we are guaranteed that a row-constrained $UE(s^{2})$-optimal design will have tight row constraints.
 In this section we investigate this question empirically.
@@
 For $c<k/2$, however, we may see tight row constraints and, as $c$ gets smaller, an increasingly large difference between the row-constrained optimal $UE(s^{2})$ value and that of the unconstrained optimal design.
-That is, if $U^{*}$ is the $UE(s^{2})$ value for the unconstrained $UE(s^{2})$-optimal design and $U_{c}^{*}$ is the $UE(s^{2})$ value of the row-constrained $UE(s^{2})$-optimal design with $c$ $+1$’s per row, then $U^{*}\leq U_{c}^{*}$ if $c\leq k/2$, but as $c\rightarrow k$, $U_{c}^{*}\rightarrow U^{*}$.
+That is, if $U^{*}$ is the $UE(s^{2})$ value for the unconstrained $UE(s^{2})$-optimal design and $U_{c}^{*}$ is the $UE(s^{2})$ value of the row-constrained $UE(s^{2})$-optimal design, then $U^{*}\leq U_{c}^{*}$ if $c\leq k/2$, but as $c\rightarrow k$, $U_{c}^{*}\rightarrow U^{*}$.
 To explore the relationship between $c$ and $UE(s^{2})$, along with the effect of $c$ on the tightness of the constraints in (6), we first examine a set of approximately $UE(s^{2})$-optimal designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, each constructed using $100$ algorithm tries of the algorithm described in Section 2.1.
-We consider several properties of the designs, each as a function of $c$: the $\text{UE}(s^{2})$ criterion value, along with the minimum, mean, and maximum slack in the row constraint, measured for each row as $(2c-k)-\sum_{j=1}^{k}x_{ij}$.
-We also plot the bounds from Section 3.1 for those designs without row constraint slack, and we provide some simulation results for various effect sizes.
-For the simulations, we generate $1{,}000$ datasets from model (1) with $\sigma^{2}=1$, $a=1$ nonzero element of $\bm{\beta}$ with value $\beta$, in line with application-specific sparsity levels of 1% or less.
-Since in these applications we know the desired effect directions, we consider $\beta\in\{0.375,0.5,0.75,1\}$, which translate to mean response differences (when the active compound is present versus absent) of $0.75$, $1$, $1.5$, and $2$, respectively.
-The datasets are analyzed using the Lasso as in (2).
-To be precise, we center $\mathbf{y}$ as $\mathbf{y}_{c}$ while also centering and scaling $X$ as $X_{cs}$ so that each column has mean 0 and the same length.
-The resulting Lasso estimates are denoted $\hat{\bm{\beta}}_{cs}$ with regularization parameter $\lambda_{cs}$.
-To reduce false positive rates, we perform a thresholding procedure in the following way.
-We assume known $\sigma^{2}$ (see discussion in Section 6 for a relaxation of this assumption) and, after obtaining $\hat{\bm{\beta}}$ for all $\lambda_{cs}$, we set any elements of $\hat{\bm{\beta}}_{cs}$ to 0 if they are less than $\sigma/8$, a value was chosen via testing to control FPR.
-For the resulting set of nonzero Lasso estimates, we obtain ordinary least squares estimates and choose the $\lambda_{cs}$ to minimize the Bayesian Information Criterion.
-The resulting set of nonzero estimates of $\bm{\beta}_{cs}$ constitute the set of factors deemed potential hits.
-The true positive rate (TPR) is the proportion of truly active factors that are correctly identified, and the false positive rate (FPR) is the proportion of truly inactive factors that are incorrectly deemed to be hits.
+We consider several properties of the designs, each as a function of $c$: the $\text{UE}(s^{2})$ criterion value, along with the minimum, mean, and maximum slack in the row constraint, where slack is measured for each row $i$ as $(2c-k)-\sum_{j=1}^{k}x_{ij}$.
+We also plot the bounds from Section 3.1 for those designs without row constraint slack, and provide some simulation results.
+For the simulations, we follow the approach we describe in Section 4, assuming $a=1$ nonzero element of $\bm{\beta}$ with value $\beta$, in line with application-specific sparsity levels of 1% or less.
+Since we know the desired effect directions, we consider $\beta\in\{0.375,0.5,0.75,1\}$, which translate to differences in mean response (pools with vs.
+pool without active compound) of $0.75$, $1$, $1.5$, and $2$, respectively.
+The datasets are analyzed using the Lasso as in Section 4.1.3.
 Figure 1 provides a visualization of $\sqrt{UE(s^{2})}$ and the measures of row-constraint slack for all $c$ (left) and for a smaller set of $c$ (right).
@@
 Thus, we conclude that for these designs any row constraint larger than $k/2=72$ will not have an appreciable effect on the results of an experiment using them, though the constraint includes slack beginning at some value $c$ slightly smaller than $k/2$.
-We have seen this same pattern for a number of other design sets as well (see Supplementary Material).
+We have seen the same pattern for a number of other design sets as well (see Section 3 of Supplementary Material A).
 The simulation (Figure 2) also shows that even for $c$ as small as 5 or 10, the designs reliably categorize active and inactive factors for sufficiently large effect sizes.
@@
 Plots are a function of $c$, the number of $+1$’s allowed in each row of design.
-Top left: $UE(s^{2})$ value for each design (blue) along with the bounds from Section 3.1 (for all designs with no row-constraint slack).
+Top left: $UE(s^{2})$ value for each design (solid line) along with the bounds from Section 3.1 (for all designs with no row-constraint slack) (dashed line).
 Bottom left: $(2c-k)-\sum_{j=1}^{k}x_{ij}$; the solid line is the average slack, while the lower and upper dashed lines are the minimum and maximum slack.
@@
 Bottom right: The same as the bottom left plot, for $40\leq c\leq 84$.
-Figure 2: Simulation results for CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, for four effect sizes, assuming one active factor; the solid lines are true positive rates while the dashed lines are false positive rates.
+Figure 2: Simulation results for CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, for four effect sizes, assuming one active factor; the heavier lines above are true positive rates while the lighter lines below are false positive rates.
 4 Comparison of Biological Screening Methods
@@
 Third, we assume a known background mean response, $\mu$, when there are no hits in a particular well.
-(We provide an alternative set of simulations in the Supplementary Material that relaxes the assumptions about $\mu$ and $\sigma^{2}$.) Finally, we assume that responses, or some standard transformation of them, can be reasonably approximated by a normal distribution.
+(We provide an alternative set of simulations in Section 5 of Supplementary Material A that relaxes the assumptions about $\mu$ and $\sigma^{2}$.) Finally, we assume that responses, or some standard transformation of them, can be reasonably approximated by a normal distribution.
 4.1.1 One-Compound-One-Well (OCOW)
@@
 4.1.2 poolHiTS
-The poolHiTS procedure, using Shifted Transversal Designs (STD, Kainkaryam and Woolf, 2008), pools compounds as we suggest and also provides for row constraints.
-However, while its design and analysis approaches are intuitive, the STD is neither constructed using a statistically motivated criterion nor analyzed using a statistical model.
-In addition to this, they suffer from the following additional disadvantages compared to our proposed procedure.
-First, due to their cyclic construction, STDs are not available for many combinations of the numbers of compounds screened, wells and compounds per well.
-For example, to screen $k=96$ compounds with $n=96$ wells, one cannot construct an STD design with $c=30$ compounds per well.
-To obtain a higher $c$, one must decrease $n$.
+The poolHiTS procedure, using Shifted Transversal Designs (STD, Kainkaryam and Woolf, 2008), is a pooling method which also provides for row constraints.
+Its designs are intuitive and easy to construct, but not based upon a statistically motivated criterion.
+Similarly, its analysis procedure is not based upon the estimation of a statistical model.
+These designs also are unavailable for many $(n,k,c)$ combinations.
 In our comparisons in Section 4, we had to choose $(n,k,c)$ values that STD allowed, rather than the most natural values.
-Real experimental screens will invariably have unique requirements, which limits the usefulness of poolHiTS.
-Second, the poolHiTS methodology assumes a binary response vector (that is, each well can be categorized as a “hit” if it contains an active compound, or a “miss” if it doesn’t), so in the standard numeric-response setting it is required to determine a threshold beyond which a well’s response is categorized as a “hit”.
-Details can be found in Kainkaryam and Woolf (2008), but here we briefly describe the basic poolHiTS methodology.
-The user specifies the total number of compounds $k$ and the maximum number of compounds per well $c$, along with the anticipated maximum number of active compounds and an overall error rate.
-This allows the construction of a design with $a$ layers, each layer consisting of $q$ pools, where the parameters $a$ and $q$ are determined by the procedure.
-This results in a final STD with $n=aq$ wells, which highlights its lack of flexibility to choose the number of wells for a given $k$ and $c$.
-The decoding algorithm requires a binary response vector; thus, to account for uncertainty associated with such a binarization, we specify a threshold, much like that for OCOW, which we can use to label each well.
-We conducted some tests to establish a threshold which would best balance TPR and FPR.
-In the end, assuming a positive effect direction, we chose $\mu+F^{-1}(0.96)\sigma$, where $F^{-1}$ is again the inverse standard normal distribution function; that is, any wells whose response fell beyond the threshold were labeled “hits” and others assumed inert.
-Once the pools have been labeled, the decoding algorithm proceeds by eliminating compounds if they appear in at least $E+1$ inert wells, and tagging compounds as active if they aren’t labeled as inert and appear in at least $E+1$ positive pools.
-Here $E$ is the number of errors allowed in the analysis.
-We note that it is possible for some compounds to fall into a third inconclusive category, and in this case we label these as active since further testing would be required to eliminate them.
+This methodology also assumes a binary response vector (that is, that each well can be categorized as a “hit” if it contains an active compound, or a “miss” if it doesn’t), so in the standard numeric-response setting it is required to determine a threshold beyond which a well’s response is categorized as a “hit”.
+We conducted some tests to establish an appropriate threshold which would best balance TPR and FPR.
+In the end, assuming a positive effect direction, we chose $\mu+F^{-1}(0.96)\sigma$, where $F^{-1}$ is again the inverse standard normal distribution function.
+A description of this method is provided in Section 4 of Supplementary Material A.
 4.1.3 Constrained Row Screening
@@
 For the simulations, where using a profile plot to choose active effects is impractical, we use a thresholding procedure to reduce FPR, with the threshold chosen via informal testing.
-In particular, our analysis procedure is:
-  • Center and scale the design $X$ so that each column has the same length, call it $X_{cs}$; and center response $\mathbf{y}$, call it $\mathbf{y}_{c}$;
+In particular, our analysis procedure is the following:
+  1. (a)
+    Center and scale the design $X$ so that each column has the same length, call it $X_{cs}$; and center response $\mathbf{y}$, call it $\mathbf{y}_{c}$.
+  2. (b)
     Obtain Lasso solutions for $\log(\lambda)$ ranging from $-8$ to $\log(\max(|X_{cs}^{T}\mathbf{y}_{c}|))$, the range chosen to reasonably ensure a full exploration of the Lasso estimation profile.
+  3. (c)
     Across all $\lambda$, set any parameter estimate to 0 if it is smaller than $\sigma/8$. Because we assume a known effect direction, we threshold all negative estimates.
+  4. (d)
     Refit each model, across $\lambda$’s, using least squares, and choose the model with the smallest BIC. The factors with nonzero estimates for this model are declared hits.
@@
 The other three experiments were more difficult in various ways, whether a smaller n/k ratio (Experiment 2) or row constraints (Experiments 3 and 4).
-The designs for each experiment were constructed using the procedure described in Section 2; they are available in the Supplementary Material.
+The designs for each experiment were constructed using the procedure described in Section 2.
 The compounds were arrayed into normalized pools implied by the designs such that each compound was at a concentration of 30 $\mu$M.
@@
 Enzymatic assays were carried out in microplates at 25C and results were determined by reading the absorbance at 495 nm in a Molecular Devices SpectraMax iD5 plate reader.
-The list of compounds, the four designs, and the experimental results are provided in the Supplementary Material.
+Additional details are in Supplementary Material A, while the list of compounds, the four designs, and the experimental results are provided in Supplementary Material D.
 Figure 4 shows the Lasso estimate profile plots for the four experiments.
@@
 6 Discussion
-In this paper, we have developed a new class of supersaturated designs, motivated by applications in high-throughput screening which use multiwell plates to search for promising chemical compounds.
-These designs constrain the number of compounds included in each well based on logistical and biological concerns of the specific HTS setting.
-These designs are used as part of a method we call CRowS that use both a statistical criterion—the $UE(s^{2})$-criterion—to guide construction of the design and a statistical analysis method—the Lasso—to obtain the set of possible hit compounds.
-Thus, our work provides a novel and improved solution to a scientific problem as well as innovative statistical design methodology inspired by the problem.
-In addition, we have demonstrated an important application of supersaturated designs, including real experimental results that display the proposed methodology in action.
-Finally, though the proof-of-concept experiments were small, we show via simulation the use of larger SSDs than have previously been typically studied, and our methodology can scale to the hundreds or thousands of factors, opening up a rich set of computational and statistical problems for the statistical community to explore.
-More specifically, in this work we have developed an efficient computational procedure to construct CRowS designs, which minimize the $UE(s^{2})$ criterion subject to $c$, the number of compounds allowed per well.
-We have also studied the behavior of these designs as a function of $c$, both theoretically and empirically.
-Theoretically, we provide initial lower bounds on the $UE(s^{2})$ criterion which can be used in some cases to verify that the globally optimal design has been obtained.
-Empirically, we find that when $c$ is between 40% and 45% of $k$, the constrained design is not appreciably worse than the unconstrained design.
-Furthermore, even severely constrained designs with $c$ as small as 5 or 10 still reliably find hit compounds in simulations.
-We also find that compared to the standard one-compound-one-well practice, or a non-statistical method that allows the pooling of multiple compounds in a well, CRowS detects hit compounds with substantially more reliability, while controlling the rate of false discoveries.
-Under the settings considered, our approach represents a clear improvement over existing HTS design and analysis methodology.
-However, there are several limitations to our work as well.
-First, though our design construction procedure is fast for the sizes of designs reported in this paper (on the order of seconds) and also for designs with $n$ and $k$ both in the hundreds (on the order of minutes), when $n$ and $k$ are both in the thousands the procedure takes on the order of days to produce a locally optimal design.
-This may not be a major concern because even a screening campaign studying thousands of compounds could be broken down into designs built on 96- or 384-well plates.
-If, however, screeners wished to experiment on, say, 1,536-well plates, they would need to be willing to wait several days for a design.
-Related to this are liquid handling challenges.
-One advantage of one-compound-one-well is that only a limited amount of liquid must be transferred from source to target plates.
-If instead of a single compound per well, we have $c$ compounds per well, then the time it takes to prepare the plates will increase proportionally to $c$.
+In this paper, we have developed and studied a new class of supersaturated designs, which we call CRowS designs, motivated by applications in high-throughput screening and in particular to the search for enzymes that cause antibiotic resistance.
+In these designs, the number of compounds per well is constrained and we use the $UE(s^{2})$ criterion to guide their construction.
+We present an efficient exchange algorithm to construct them, present some initial lower bounds on the criterion value, study the effect of the row constraint upon the designs, show that they compare quite favorably to common competitors, and demonstrate that they detect a metallo-$\beta$-lactamase inhibitor in real experiments.
+Thus, our work presents an innovative statistical design methodology inspired by an important scientific problem, along with the suggestion to use a statistical modeling approach to the analysis of the designs.
+In addition, these designs are larger than have typically been studied, and our methodology can scale to the hundreds or thousands of factors.
+Though our approach represents a clear improvement over traditional HTS methodology, it also has several limitations.
+First, while our design construction procedure is fast for the sizes of designs reported in this paper, for designs with $n$ and $k$ both in the thousands (e.g.
+an experiment on a 1,536-well plate) may require on the order of days to produce an approximately optimal design.
+Also, this methodology should save on the number of required plates compared to one-compound-one-well, but may require more liquid handling since each well includes $c$ different compounds.
 This, of course, must be weighed against the improved hit rates that pooling using our method could provide.
-Another important concern is interactions.
-In this work, we have assumed that compounds don’t interact either statistically or chemically.
-A natural question is whether statistical interactions between factors might affect the CRowS method’s ability to detect hits.
-In order to investigate this we repeated the simulations described in Section 4 where the response was generated using the model in (1) but the $\bm{\beta}$ included two active factors and a single active interaction.
+We also note that depending on the application, statistical or chemical interactions could be a concern.
+We briefly investigated a setting in which statistical interactions are sparse (this could arise, for instance, if interactions only occur between compounds that are both hits), by repeating the simulations described in Section 4 where the response was generated using the model in (2) but the $\bm{\beta}$ included two active factors and a single active interaction, half the size of the active main effects.
 We considered both synergistic and antagonistic interactions that obey strong (interaction only appears when both parents are active) or weak (interaction only includes one active parent) heredity.
-In the Supplementary Material, we provide results of this simulation.
-We find that synergistic interactions degrade the ability of CRowS to detect hit compounds, while antagonistic interactions improve the results.
-We do note that Singh and Stufken (2023) has recently proposed a method to estimate interactions in SSD’s.
-We are currently working to adapt and apply Singh and Stufken (2023) to the present setting, so that important interactions can be accounted for and/or estimated.
-On the application side, our work leads to a number of interesting questions.
-Though there are numerous examples of successful pooling in the literature (e.g.
-Devlin et al., 1996; Wilson-Lingardo et al., 1996; Taylor et al., 1997) and even in the context of commercial drug discovery (Elkin et al., 2015), there is also a concern regarding so-called “promiscuous interactions” (Feng and Shoichet, 2006), in which compounds may react to each other in unpredictable ways.
-In our own experimental work, before fine-tuning the process by which the designs were implemented we twice failed to achieve successful results.
-In the first attempt in particular, we were experimenting on a wider set of compounds than the group we have reported on in this paper.
-Additional experiments may clarify from the chemical side the limits of pooling in this or other contexts.
-This work raises a number of theoretical questions as well.
-From our empirical work, it appears that row-constraint slack begins to appear at between 40% and 45% of $k$.
-Can this be characterized more precisely?
-Can we specify a threshold for $c$ beyond which we are guaranteed that we will have no row-constraint slack?
-Can we tighten the bounds on the designs, and perhaps develop construction methods for certain $(n,k,c)$ combinations that bypass the need for an algorithm altogether?
+In Section 7 of Supplementary Material A, we find that synergistic interactions degrade the ability of CRowS to detect hit compounds, while antagonistic interactions improve the results.
+We note that Singh and Stufken (2023) has recently proposed a method to estimate interactions in SSD’s, and this could provide a possible solution to the problem of sparse interactions.
+There is also the possibility of “promiscuous inhibitors” (Feng and Shoichet, 2006), in which compounds may react to each other in unpredictable ways.
+This problem may be application-specific, or specific to particular chemical groups, and warrants additional exploration in future studies seeking to expand the number of compounds that can be simultaneously tested.
+Our work raises a number of interesting follow-up questions as well.
+On the application side, can the methodology be used in larger-scale settings to identify potential hits?
+In large-scale screens, can our proposed methodology be adjusted to largely eliminate false positives in large effect environments, or be used as a secondary screening tool to eliminate false positives that are identified in the initial screen?
+Theoretically, can the lower bounds be tightened and the reliance on the assumption of tight constraints be relaxed?
+More fundamentally, is it possible to more rigorously characterize conditions that will guarantee tight row-constraints, rather than the empirical observation we’ve made that row-constraints are tight until $c$ is larger than 40%-45% of $k$?
+We leave these questions for future work.
+In general, our work opens up new applications and development for SSDs.
+These CRowS designs can be applied to a broad class of biological and chemical screening problems, and the size of the designs are much larger than those that have been typically studied in the SSD literature.
+We note that Eskridge et al.
+(2004) study large SSDs, and even give a nod to HTS as an application, but they don’t provide designs for $n>24$ nor do they consider row constraints or provide a real example.
 Supplementary Material
@@
     Designs, data, and analysis of the real experiments.
-Acknowledgements, Funding, and Disclosure Statement
+Acknowledgements
 This work was funded in part by Miami University’s Faculty Research Committee (BJS) and by the National Institutes of Health through grant GM128595 (RCP).
 We also thank Jens Mueller from Miami’s Research Computing Support for his computational assistance.
-The authors report there are no competing interests to declare.
+Disclosure and Data Availability Statement
+The authors report there are no competing interests to declare. The data underlying this article are available in its online supplementary material.